OpenStudy (anonymous):

6m-5=7m+7-m soloution or no sloution?

OpenStudy (agreene):

no solution. 6m-5=7m+7-m combine likes on the right 6m-5=6m+7 subtract 6m from both sides and add 5 0=12 <== not true.

OpenStudy (mathteacher1729):

Or you could write it like this: 6m - 5 = 6m +7 same slope (6), different y-intercepts (-5) and (+7). That means they are parallel, but not overlapping, and there is no solution. Picture attached.